find all the solutions on the interval [0,2pi).
2tan2x-2sqrt3=0

Answer by stanbon(75887)

find all the solutions on the interval [0,2pi).
2tan2x-2sqrt3=0
----
tan(2x) - sqrt(3) = 0
----
tan(2x) = sqrt(3) = 0
2x = tan^-1[sqrt(3)]
---
2x = pi/3 or (4/3)pi
---
x = pi/6 or (2/3)pi
----
Cheers,
Stan H.
=========================